OpenStudy (sonyalee77):

Write the equation of the line that passes through( -4,3)and is perpendicular to the line y=7/5x-3/5

OpenStudy (anonymous):

Two perpendicular lines have slopes which are "flipped" and have different signs. For example, lets say a line had a slope of (3/2). A perpendicular line would have a slope of (-2/3)

OpenStudy (anonymous):

Other than that, you do the same thing as before. Is the equation y=7/((5x-3)/5) (Does it have a complex fraction?

OpenStudy (anonymous):

Is this the equation? You don't use ( or ) so I want to make sure

OpenStudy (anonymous):

Or is this the equation:

OpenStudy (anonymous):

Can you draw out the equation as it appears? I can't help you, since the two equations are very different from each other

OpenStudy (anonymous):

*unless I know which one it is

OpenStudy (sonyalee77):

A.5x+7y=1 B.7x+5y=-13 C.7x+5y=-1 D.3x+5y=7

OpenStudy (anonymous):

I keep getting the wrong answer because I am going off the equation y = (7/5)x -(3/5)

OpenStudy (sonyalee77):

@superdavesuper

OpenStudy (anonymous):

So have a line that is perpendicular the slopes have to be opposite So the slope of the other equation for be (-5/7) Now that we know that we start with y-y1=m(x-x1) Plug in what we know ( -4,3) ((remember (x,y)), and m=(-5/7) So y-(3)=(-5/7)(x+4) y-3=-5/7x-23/7 Add 3 to both sides y=-5/7x-3/7 Multiply both sides by 7 7y=-5x-3 Add 5x to both sides 5x+7y=-3 This is what I got

OpenStudy (triciaal):

y = 7/5 x -3/5 perpendicular line so slope is -1/m =-5/7 point on line (-4,3) 3 = -5/7 (-4) + b b = 1/7 eqn y = -5/7 x + 1/7 multiply by 7 to get rid of fraction rewrite to match format of answer 7y +5x = 1

OpenStudy (triciaal):

rosedryer had the right idea but this line y-(3)=(-5/7)(x+4) had an error x = -4 need to find b